How can we write log n (v / v0) = 2x as v = v0e2x ? Please explain.

Grade:12

2 Answers

Godfrey Classic Prince
12 years ago

Dear Manish Sahu,

This is a rule of logarithm that 1/ln (x) becomes ex .

This is the same thing that has happened here .

Logn has gone as a denominator to the other side making it exponential !!

 

Hope this helped you immensely..

All the Very Best & Good Luck to you ...

Regards,

AskIITians Expert,

Godfrey Classic Prince

IIT-Madras

Please do not forget to approve my answer if you liked it by clicking on "Yes" given below...!!Smile

Riddheish Seth
12 years ago

log v - log vo = 2x (log to base n)

lnv - lnvo = 2x ln (n)

ln (v/vo) = 2x ln(n)

e^2xln(n) = v/vo